Brown Bank is designing a new account that pays interest quarterly. They want to pay effectively, a 16% per year on this account. They want to advertise the APR on this new account. What is the Apr that corresponds to an effective rate of 16% for this new account? Explain.
Answer #1

EAR=(1+APR/m)^m-1
where m=compounding periods

0.16=(1+APR/4)^4-1

(1+0.16)^(1/4)=(1+APR/4)

APR=[(1+0.16)^(1/4)-1]*4

= 15.12%(Approx).